The pH of a solution of HBr is found to be 3. How many grams of HBr are in a 100 mL sample?
Answer #1

use:

pH = -log [H+]

3 = -log [H+]

[H+] = 1*10^-3 M

volume , V = 1*10^2 mL

= 0.1 L

use:

number of mol,

n = Molarity * Volume

= 1*10^-3*0.1

= 1*10^-4 mol

Molar mass of HBr,

MM = 1*MM(H) + 1*MM(Br)

= 1*1.008 + 1*79.9

= 80.908 g/mol

use:

mass of HBr,

m = number of mol * molar mass

= 1*10^-4 mol * 80.91 g/mol

= 8.091*10^-3 g

Answer: 8.09*10^-3 g